# Decoherence and Probability

### v2

@@ 3.3.1. Classical Possibility Space Models
The conditions on the representation VII–X encode two features which will be important for the comparison with phase space representations of quantum possibly spaces. These are the conservation and localisability of probability density.
The conservation of probability density is a well known feature of a phase space representations of a classical possibly model. It is typically expressed via the Liouville equation:
This equation guarantees that the additivity property of regions of phase space is preserved over time. If we think of the probability like a fluid, we can understand there to be a phase space 3-current given by the tuple . The Liouville equation then expresses the conservation of the 3-current and is equivalent to the statement that for any region the net ‘efflux’ of probability is zero. This is the characteristic property of a fluid with an incompressible flow and is a result of the absence sources or sinks of the probability ‘fluid’ (Pathria and Beale 2011, p. 28), cf. (Gibbs 1902, p.11).
@@
A stochastic phase space model provides a partial interpretation of a classical probabilistic structure. The state space is the sample space . The Borel sets given by sub-regions of phase space are the -algebra (Feller 1991). The probability measure is given by the integration of the probability density function with respect to the volume measure over a sub-region . The connection between the conditions IXXXiii and VIVIiii is guaranteed by the definition of -algebra. The event corresponding to the entire space is certain. The event corresponding to the empty set is impossible. The probability measure provides us a representations of events being possible and being higher weighted. The model includes a deterministic sub-set since a function that approximates a -function is an admissible PDF and thus the case in which the singleton of the Borel sets is measure one and all other points are measure zero is an admissible stochastic phase space model. It is worth emphasising that a stochastic phase space model is not a full interpretation of a classical probabilistic structure since the weighting of possibility provided in terms of the probability measure admits further admissible interpretation in terms of standard notions of credence, frequency, or physical chance.
@@ 3.3.2. Quantum Possibility Space Models
The contrast between classical and quantum possibly space models is greatly clarified by examining the failure of conservation and localisability of quasi-probability density implied by the conditions XVIII–XXI. Let us demonstrate this failure explicitly for the choice of Wigner function, as the quasi-probability density function.666The Wigner function is the most important of the quasi-probability distributions on quantum phase space that can be defined via different operator ordering conventions. Our discussion principally draws upon details in Curtright et al. (2013) unless otherwise noted. See Section 4 for further details on the Wigner function.
@@
Failure of conservation of quasi-probability is a direct consequence of the non-commutativity of the Moyal algebra of quantum phase space observables in comparison to the Poisson algebra of classical phase space observables as encoded in the relation . We can show this explicitly by considering the quasi-probability flux for some arbitrary region of phase space with volume greater than or equal to the minimum volume. This is given by the expression (Curtright et al. 2013, p. 57):
@@
where we have used the Wigner transform of the Heisenberg equations of motion and and the Moyal equation . The quasi-probability density associated with regions of phase space thus manifests a violation of additivity over time in marked contrast to the classical probability density function in phase space, cf. (Wallace 2021, p.23).
@@
The failure of localisability can be understood as follows. Although the quasi-probability functions need not in general be Lebesgue integrable over the entire phase space (Aniello 2016) one can show that, for example, the Wigner function is an element of and is thus a square integrable and continuous function on the phase space (Landsman 2012, p. 142). Moreover, the induced quasi-measure is -additive and the Wigner function induces a finite signed measure (Dias et al. 2019). It might seem therefore that besides the conservation failure and the negativity the Wigner function is a very much like a classical probability function. The bound, however, leads to an important difference in the types of possibility space representations that a quasi-probability function can provide. By the Cauchy–Schwarz inequality the function is bounded such that and we thus have that . This, in turn, leads to a restriction of to volumes of phase space greater than equal to one in units of (Dell’Antonio 2016, p.19). Thus, in contrast to the classical case, it is not possible to concentrate quasi-probability density almost entirely at a single point. This amounts to precluding the possibility that the quasi-probability density function approximates a -function in phase space (Leonhardt 2010, p.71). Phase space points are not in and we cannot have a situation in which the Wigner function is non-zero at a point but zero (almost) everywhere else. Physically speaking, non-localisablity can be understood as a consequence of the Heisenberg uncertainty principle which, in turn, is a direct consequence of the non-commutative structure induced by the -product. See (Curtright et al. 2013, §5) and (Huggett et al. 2021, §5.1). The structure of a quantum possibility space model encodes the fact that quantum possibilities are not ‘distinct’ in the sense that the failure of localisability implies we can only ‘peak’ the Wigner function on regions of finite size to a limited extent. This is in addition to the failure of ‘distinctness’ of possibilities via the existence of entanglement which is encoded in the negativity of the Wigner function (Dahl et al. 2006).
@@ 4. Probability and Classicality
In this section we demonstrate that the combination of ‘quasi-probabilistic emergence’ and ‘semi-classical emergence’ allow us to derive a classical possibility space model from a quantum possibility space model. This demonstration depends upon two important results. First, that explicit models of decoherence in quantum phase space display the generic feature that the Wigner quasi-probability distribution is positive-definite after finite times of the order of the decoherence time. This is the quasi-probabilistic emergence with Wigner positivity the relevant novel and robust behaviour. Second, that the generalised Ehrenfest relations imply that the classical and quantum moment evolution equations are syntactically isomorphic with the Wigner function playing the role of a probability density function. A positive-definite Wigner function then displays localisation and conservation behaviour identical to that of a probability density function to the extent to which we can neglect terms . This is the semi-classical emergence with localisation and conservation the relevant novel and robust behaviour.
@@ 4.1. Wigner Negativity and Decoherence
The Wigner function is the central object in the phase space approach to quantum mechanics.777A concise and very clear introduction to the Wigner function and the quantum phase space formalism is Curtright et al. (2013). Further useful discussions can be found in O’Connell and Wigner (1981); Hillery et al. (1984); Case (2008); De Gosson (2017); Leonhardt (2010). The small philosophical literature is principally comprised of the discussions found in Suppes (1961); Cohen (1966); Sneed (1970); Friederich (2021); Wallace (2021). Representing the quantum state of a system via a density matrix, , the Wigner function, , takes the form:
@@
Significantly, it can be proved that any quasi-probability distribution function of the form which and reproduces the marginal probability densities cannot also be positive semi-definite (Wigner 1971). The Wigner function thus cannot be positive semi-definite. Wigner negativity has been variously recognised as the distinctive non-classical feature of the Wigner function and has been shown to have direct implications for both contextually and entanglement (Dahl et al. 2006; Delfosse et al. 2017; Booth et al. 2022). The size of the regions of negativity in phase space are of order which will be important in what follows. Significantly, the sub-set of Wigner functions that correspond to minimum uncertainty coherent states can be shown to be everywhere positive (and visa versa) (Hudson 1974; Mariño 2021).
Despite its negativity, the Wigner function has a number of attractive features that mark it out as privileged among the quasi-probability distribution functions. In particular, the density and marginal features noted above crucially depend upon the -product associated to the Wigner function being the Moyal -product. This is what allows one -product to be dropped inside an integral via integration by parts leading to formal behaviour that matches that of a genuine probability density function for the marginals and expectation values.888This feature is in contrast to the Husimi Q-function for which the associated -product cannot be integrated out and leads to marginals distributions that do not correspond to those of quantum mechanics (Curtright et al. 2013, §13). For further discussion of quasi-probability distributions and probability interpretations see Leonhardt (2010); Schroeck (2013); Friederich (2021); Stoica (2021); Umekawa et al. (2024).
Let us now consider the behaviour of the Wigner function within a simple model of decoherence with a focus on the role of Wigner negativity. The general framework for the study of decoherence is quantum master equations for the reduced density matrix of a quantum system. For our purposes it will suffice to consider the most basic master equation, that due to Joos and Zeh (1985). The Joos-Zeh equation can be derived based on a idealised decoherence model with recoilless scattering that carries away information but not momentum of a quantum particle. It is a minimal model for position localisation of a quantum particle via the destruction of coherence. More realistic models include noise and dissipation terms but share the central formal feature of Gaussian-smoothing.
@@
Following Diósi and Kiefer (2002), the Fokker-Plank equation for the Wigner function can be demonstrated to be equivalent to a progressive Gaussian-smoothing of an initial Wigner function . In particular, we can re-write the Equation (10) as a convolution of the form:
where is a generalised Gaussian function with time dependent correlation matrix:
@@
Convolution with a Gaussian function, as per the heat equation, has the general effect of smoothing the Wigner function.999More generally, we can understand decoherence in terms of convolution of the Wigner function with a Gaussian according to a Weierstrass transform. This is, in fact, precisely to transform a Wigner function into a Husimi Q-function (Curtright et al. 2013, §13). We should not expect the quantum mechanical marginal probabilities to be fully recoverable from the reduced state post-decoherence. Which is perhaps unsurprising. The regions of Wigner negativity are of order and a Gaussian smoothing can be shown to be such that it will progressively render any initial Wigner function positive-definite.101010This is true for Gaussian smoothings but does not hold in general for any averaging (de Aguiar and de Almeida 1990). Indeed, Diósi and Kiefer (2002) show that by Equation (11), any initial state will be such that Wigner function will be strictly positive after a finite time which is of the order of the decoherence timescale defined above. The result of Diósi and Kiefer (2002) demonstrates that even for the most simple model of decoherence the dynamical equations serve to smooth out structure of the Wigner function and eliminate Wigner negativity almost immediately.111111See (Brody et al. 2024) for a quantum dynamical model that achieves Wigner positivity in finite time without a von Neumann term, i.e. the first commutator on the right hand side of (9) in the Joos-Zeh-model, but with a different, but still simple, form of the dissipator term. Generically, we can expect that Wigner positivity is a novel and robust behaviour that emerges via decoherence based upon a partially interpreted quasi-probability structure.
The generic dynamical phenomena of Wigner positivity illustrates the importance of differentiating between classical probability structures and classical possibility space models. The measure induced by a positive Wigner function are realisations of classical probability structures satisfying the Kolmogorov axioms. However, such a measure cannot be interpreted in terms of a classical possibility space models. A positive Wigner function is not equivalent to a classical probability density function since the crucial features of conservation and localisation will still fail, notwithstanding the Wigner function being positive. On appropriate scales, we will still find the Gaussian-smoothed, positive Wigner function acting in a manner that is irreconcilable with it being a classical probability density and thus, as it stands, we cannot apply an interpretation in terms of a classical possibility space model. To our knowledge this simple but extremely important point has not previously been highlighted in the literature.
@@ 4.2. The Role of Semi-Classicality
The previous section provided a simple illustration of how the non-classical feature of Wigner negativity can be eliminated via decoherence. These methods as a basis for describing the emergence of classicality can generalised to more realistic models.
Famously, the approach was extended to the study of non-linear models, such as that of the classically chaotic orbit of Hyperion, by Habib et al. (1998) leading to the iconic illustrations reproduced in Figure 1. The figures (a) and (b) both show the Wigner distribution function from a solution a Fokker-Planck type equation for a non-linear system with a quartic term in the Hamiltonian. The difference between the figures corresponds to solutions to the model at a given time without (a) and with (b) the destruction of large scale quantum coherence. The figure (c) shows the solution of a classical Fokker-Planck equation for a classical probability density function. The box represents a phase space area of .
The model of Habib et al. (1998) provides a vivid exemplification of what it means for a quantum possibility space model evolving under an open quantum dynamics to exhibit behaviour that displays a remarkable close correspondence to that exhibited by a classical possibility space model. In a rich and insightful analysis of the model, Franklin (2023) proposes that we can use appeal to decoherence to frame an account of the emergence classical chaotic phenomenology based upon an underlying non-chaotic quantum state.121212Franklin’s account builds upon on the account of emergence described in Franklin and Robertson (2021) which in turn builds on Ross (2000). See also Ladyman and Ross (2007), Wallace (2010) and Mulder (2024). The claim runs as follows:
@@
There is much to recommend in Franklin’s analysis as an account of the relationship between classical and quantum phenomenology. However, one must also bear in mind the foregoing detailed treatment of the structure of decoherence models based upon the Wigner function. Evidently, a partial interpretation of a quasi-probability structure via possibility space model is a necessary ingredient within any model of decoherence based upon the Wigner function (or, arguably, more generally). One is already making use of a partially interpreted generalised probabilistic structure – a possibility weighting – when one models decoherence via dynamical equations for the Wigner function as derived from quantum master equations.
Classical possibility space models do not emerge ab initio from a non-probabilistic and uninterpreted formalism, but rather are emergent in the relevant sense from a partially interpreted, quasi-probabilistic structure. The model of the emergence of classical phenomenology that decoherence models based upon the Wigner function provide is explicitly reliant on its role as a quasi-probability density function in inducing a possibility space weighting. Decoherence must be understood as a basis for quasi-probabilistic emergence rather than non-probabilistic emergence. It is also worth noting that the justification of the Wigner function as the privileged representation of quasi-probability relies upon its unique ability to recover the experimentally confirmed marginals given by Born rule probabilities.
Furthermore, as already noted, Wigner positivity is necessary but not sufficient for us to interpret a model as a representation of a classical possibility space due to the failure of localisability and conservation. To understand the relation between the two one is required to provide an account of the semi-classical limiting relation between the two which we will in the following section. What we can show, however, is that there is an isomorphism between the dynamical moment equations in quantum and classical possibly space formalisms with the Wigner function and probability density function playing identical roles in the relevant equations, cf. (Wallace 2021, p.23).
@@
Following Ballentine and McRae (1998), the corresponding formulas in the classical possibility space model is:
We thus have an isomorphism between the classical and quantum probabilistic phase space formalism. The classical and quantum moment evolution first equations are syntactically isomorphic with the Wigner function playing the role of a probability density function.

### v3

@@ 3.3.1. Classical Possibility Space Models
The conditions on the representation VII–X encode two features which will be important for the comparison with phase space representations of quantum possibly spaces. These are the local conservation and localisability of probability density.
The local conservation of probability density is a well known feature of a phase space representations of a classical possibly model. It is typically expressed via the Liouville equation:
This equation guarantees that the additivity property of regions of phase space is locally preserved over time. If we think of the probability like a fluid, we can understand there to be a phase space 3-current given by the tuple . The Liouville equation then expresses the fact that the current represents an incompressible flow and is a result of the absence of compression or rarefaction points in the probability ‘fluid’ (Pathria and Beale 2011, p. 28), cf. (Gibbs 1902, p.11) and the density about the representative point in the flow is conserved.
@@
A stochastic phase space model provides a partial interpretation of a classical probabilistic structure. The state space is the sample space . The Borel sets given by sub-regions of phase space are the -algebra (Feller 1991). The probability measure is given by the integration of the probability density function with respect to the volume measure over a sub-region . The connection between the conditions IXXIIIiii and VIVIiii is guaranteed by the definition of -algebra. The event corresponding to the entire space is certain. The event corresponding to the empty set is impossible. The probability measure provides us a representations of events being possible and being higher weighted. The model includes a deterministic subset since a function that approximates a -function is an admissible PDF and thus the case in which the singleton of the Borel sets is measure one and all other points are measure zero is an admissible stochastic phase space model. It is worth emphasising that a stochastic phase space model is not a full interpretation of a classical probabilistic structure since the weighting of possibility provided in terms of the probability measure admits further admissible interpretation in terms of standard notions of credence, frequency, or physical chance.
@@ 3.3.2. Quantum Possibility Space Models
The contrast between classical and quantum possibly space models is greatly clarified by examining the failure of local conservation and localisability of quasi-probability density implied by the conditions XI–XIV. Let us demonstrate this failure explicitly for the choice of Wigner function, as the quasi-probability density function (Curtright et al. 2013).
@@
Failure of local conservation of quasi-probability is a direct consequence of the non-commutativity of the Moyal algebra of quantum phase space observables in comparison to the Poisson algebra of classical phase space observables as encoded in the relation . We can show this explicitly by considering the quasi-probability flux for some arbitrary region of phase space . This is given by the expression (Curtright et al. 2013, p. 57):
@@
where we have used the Wigner transform of the Heisenberg equations of motion and and the Moyal equation . The quasi-probability density associated with regions of phase space thus manifests a violation of local additivity over time in terms of failure of conservation of the probability density in the precise sense of the density about the representative point in the flow not being conserved, cf. (Curtright et al. 2013, p. 58). This is despite retaining satisfying additivity at a time in terms of -additivity. The failure of local conservation in this sense is in marked contrast to the classical probability density function in phase space, cf. (Wallace 2021, p.23).
@@
The failure of localisability can be understood as follows. Although a quasi-probability function need not in general be Lebesgue integrable over the entire phase space (Aniello 2016) one can show that, for example, the Wigner function is an element of and is thus a square integrable and continuous function on the phase space (Landsman 2012, p. 142). Moreover, the induced quasi-measure is -additive and the Wigner function induces a finite signed measure (Dias et al. 2019). It might seem therefore that besides the conservation failure and the negativity the Wigner function is a very much like a classical probability function. The bound, however, leads to an important difference in the types of possibility space representations that a quasi-probability function can provide. By the Cauchy–Schwarz inequality the function is bounded such that and we thus have that . This, in turn, leads to a restriction of to volumes of phase space greater than equal to one in units of (Dell’Antonio 2016, p.19). Thus, in contrast to the classical case, it is not possible to concentrate quasi-probability density almost entirely at a single point. This amounts to precluding the possibility that the quasi-probability density function approximates a -function in phase space (Leonhardt 2010, p.71). Phase space points are not in and we cannot have a situation in which the Wigner function is non-zero at a point but zero (almost) everywhere else. Physically speaking, non-localisablity can be understood as a consequence of the Heisenberg uncertainty principle which, in turn, is a direct consequence of the non-commutative structure induced by the -product. See (Curtright et al. 2013, §5) and (Huggett et al. 2021, §5.1). The structure of a quantum possibility space model encodes the fact that quantum possibilities are not ‘distinct’ in the sense that the failure of localisability implies we can only ‘peak’ the Wigner function on regions of finite size to a limited extent. This is in addition to the failure of ‘distinctness’ of possibilities via the existence of entanglement which is encoded in the negativity of the Wigner function (Dahl et al. 2006).
@@ 4. Probability and Classicality
In this section we demonstrate that the combination of ‘quasi-probabilistic emergence’ and ‘semi-classical emergence’ allow us to derive a classical possibility space model from a quantum possibility space model. This demonstration depends upon two important results. First, that explicit models of decoherence in quantum phase space display the generic feature that the Wigner quasi-probability distribution is positive-definite after finite times of the order of the decoherence time. This is the quasi-probabilistic emergence with Wigner positivity the relevant novel and robust behaviour. Second, that the generalised Ehrenfest relations imply that the classical and quantum moment evolution equations are syntactically isomorphic with the Wigner function with the latter playing the same role as the probability density function in the former. A positive-definite Wigner function then displays localisation and local conservation behaviour identical to that of a probability density function to the extent to which we can neglect terms . This is the semi-classical emergence with localisation and conservation of the relevant novel and robust behaviour.
@@ 4.1. Wigner Negativity and Decoherence
The Wigner function is the central object in the phase space approach to quantum mechanics.555Physics references are O’Connell and Wigner (1981); Hillery et al. (1984); Case (2008); De Gosson (2017); Leonhardt (2010); Curtright et al. (2013). See Suppes (1961); Cohen (1966); Sneed (1970); Friederich (2021); Wallace (2021) for philosophical analysis. Representing the quantum state of a system via a density matrix, , the Wigner function, , takes the form:
@@
It can be proved that any quasi-probability distribution function of the form which reproduces the marginal probability densities corresponding to the Born rule cannot also be positive semi-definite (Wigner 1971). The Wigner function thus cannot be positive semi-definite. Wigner negativity has been variously recognised as the distinctive non-classical feature of the Wigner function and has been shown to have direct implications for both contextually and entanglement (Dahl et al. 2006; Delfosse et al. 2017; Booth et al. 2022). The size of the regions of negativity in phase space are of order . The subset of Wigner functions that correspond to minimum uncertainty coherent states can be shown to be everywhere positive (and vice versa) (Hudson 1974; Mariño 2021).
Despite its negativity, the Wigner function has a number of attractive features that mark it out as privileged among the quasi-probability distribution functions. In particular, the density and marginal features noted above crucially depend upon the -product associated to the Wigner function being the Moyal -product. This is what allows one -product to be dropped inside an integral via integration by parts, leading to formal behaviour that matches that of a genuine probability density function for the marginals and expectation values. This feature is in contrast to other -product association rules and quasi-probability distributions (such as the Husimi Q-function) which do not reproduce the full set of marginal distributions corresponding to the Born rule (Curtright et al. 2013, §13).666For further discussion see Leonhardt (2010); Schroeck (2013); Friederich (2021); Stoica (2021); Umekawa et al. (2024).
Let us now consider the behaviour of the Wigner function within a simple model of decoherence with a focus on the role of Wigner negativity. The general framework for the study of decoherence is quantum master equations for the reduced density matrix of a quantum system. For our purposes it will suffice to consider the most basic master equation, that due to Joos and Zeh (1985). The Joos-Zeh equation can be derived based on an idealised decoherence model with recoilless scattering that “carries away information about the position of the particle” (Joos et al. 2013, p. 82) – that is, induces loss of informational or von Neumann entropy – but is conservative with respect to energy and momentum. It is a minimal model for position localisation of a quantum particle via the destruction of coherence. More realistic models include noise and dissipation terms but share the central formal feature of Gaussian-smoothing.
@@
Following Diósi and Kiefer (2002), the Fokker-Plank equation for the Wigner function can be demonstrated to be equivalent to a progressive Gaussian-smoothing of an initial Wigner function . In particular, we can rewrite the Equation (10) as a convolution of the form:
where is a generalised Gaussian function with time-dependent correlation matrix:
@@
Convolution with a Gaussian function, as per the heat equation, has the general effect of smoothing the Wigner function.777More generally, we can understand decoherence in terms of convolution of the Wigner function with a Gaussian according to a Weierstrass transform. This is, in fact, precisely to transform a Wigner function into a Husimi Q-function (Curtright et al. 2013, §13). We should not expect the quantum mechanical marginal probabilities to be fully recoverable from the reduced state post-decoherence. Which is perhaps unsurprising. The regions of Wigner negativity are of order and a Gaussian smoothing can be shown to be such that it will progressively render any initial Wigner function positive-definite.888This is true for Gaussian smoothings but does not hold in general for any averaging (de Aguiar and de Almeida 1990). Indeed, Diósi and Kiefer (2002) show that by Equation (11), any initial state will be such that Wigner function will be strictly positive after a finite time which is of the order of the decoherence timescale defined above. The result of Diósi and Kiefer (2002) demonstrates that even for the simplest model of decoherence the dynamical equations smooth out structure of the Wigner function and eliminate Wigner negativity almost immediately.999See (Brody et al. 2025) for a quantum dynamical model that achieves Wigner positivity in finite time without a von Neumann term, i.e. the first commutator on the right hand side of (9) in the Joos-Zeh-model, but with a different, but still simple, form of the dissipator term. Generically, we can expect that Wigner positivity is a novel and robust behaviour that emerges via decoherence based upon the partially interpreted quasi-probability structure as provided by a a quantum possibility space model.
The generic dynamical phenomena of Wigner positivity illustrates the importance of differentiating between classical probability structures and classical possibility space models. The measure induced by a positive Wigner function are realisations of classical probability structures satisfying the Kolmogorov axioms. However, such a measure cannot be interpreted in terms of a classical possibility space models. A positive Wigner function is not equivalent to a classical probability density function since the crucial features of local conservation and localisation fail, notwithstanding the Wigner function being positive. On appropriate scales, we will still find the Gaussian-smoothed, positive Wigner function acting in a manner that is irreconcilable with it being a classical probability density and thus, as it stands, we cannot apply an interpretation in terms of a classical possibility space model. To our knowledge this important point has not previously been highlighted in the literature.
@@ 4.2. The Role of Semi-Classicality
The previous sub-section provided a simple illustration of how the non-classical feature of Wigner negativity can be eliminated via decoherence. These methods as a basis for describing the emergence of classicality can generalised to more realistic models.
Famously, the approach was extended to the study of non-linear models, such as that of the classically chaotic orbit of Hyperion, by Habib et al. (1998), leading to the iconic illustrations reproduced in Figure 1. The figures (a) and (b) both show the Wigner distribution function from a solution a Fokker-Planck type equation for a non-linear system with a quartic term in the Hamiltonian. The difference between the figures corresponds to solutions to the model at a given time without (a) and with (b) the destruction of large scale quantum coherence. The figure (c) shows the solution of a classical, Fokker-Planck equation for a classical probability density function. The box represents a phase space area of .
The model of Habib et al. (1998) provides a vivid exemplification of a quantum possibility space model evolving under an open quantum dynamics that displays a remarkable close correspondence to the behaviour exhibited by a classical possibility space model. Franklin (2023) presents the claim that we can use appeal to decoherence to frame an account of the emergence classical chaotic phenomenology based upon an underlying non-chaotic quantum state in the following terms:101010See Ross (2000); Ladyman and Ross (2007); Wallace (2010); Franklin and Robertson (2021); Mulder (2024).
@@
There is much to recommend in Franklin’s analysis as an account of the relationship between classical and quantum phenomenology. However, one must also bear in mind the foregoing detailed treatment of the structure of decoherence models based upon the Wigner function. A model of decoherence based upon the Wigner function is a partial interpretation of a quasi-probability structure via possibility space model. One is already making use of a partially interpreted generalised probabilistic structure – a possibility weighting – when one models decoherence via dynamical equations for the Wigner function as derived from quantum master equations.
Purely formal derivations do not have physical content. Classical possibility space models do not emerge ab initio from a non-probabilistic and uninterpreted formalism, but rather are emergent in the relevant sense from a partially interpreted, quasi-probabilistic structure. The model of the emergence of classical phenomenology that decoherence models based upon the Wigner function provide is explicitly reliant on its role as a quasi-probability density function in inducing a possibility space weighting. Decoherence must be understood as a basis for quasi-probabilistic emergence rather than non-probabilistic emergence. It is also worth noting that the justification for treating the Wigner function as the preferred representation of quasi-probability relies upon its privileged empirical position among quasi-probability representations and corresponding -product association rules. That is, as noted in the previous sub-section, it is only the Wigner function together with the Moyal -product that leads to the full set of experimentally confirmed marginal distributions corresponding to the Born rule.
Furthermore, as already noted, Wigner positivity is necessary but not sufficient for us to interpret a model as a representation of a classical possibility space due to the failure of localisability and conservation. To understand the relation between the two one is required to provide an account of the semi-classical limiting relation between the two which we will in the following section. There is an isomorphism, however, between the dynamical moment equations in quantum and classical possibly space formalisms with the Wigner function and probability density function playing identical roles in the relevant equations, cf. (Wallace 2021, p.23).
@@
Following Ballentine and McRae (1998), the corresponding formulas in the classical possibility space model are:
We thus have an isomorphism between the classical and quantum probabilistic phase space formalism. The classical and quantum first moment evolution equations are “syntactically isomorphic” (Hempel 1965; Bartha 2010). That is, they can be obtained by assigning different physical interpretations to the symbols that appear in a common mathematical form, with the Wigner function in the latter playing the same role as the probability density function in the former.